# ReXGroundingCT: A 3D Chest CT Dataset for Segmentation of Findings from Free-Text Reports

### v1

@@ Abstract
We present ReXGroundingCT, the first publicly available dataset to link free-text radiology findings with pixel-level segmentations in 3D chest CT scans that is manually annotated. While prior datasets have relied on structured labels or predefined categories, ReXGroundingCT captures the full expressiveness of clinical language represented in free text and grounds it to spatially localized 3D segmentation annotations in volumetric imaging. This addresses a critical gap in medical AI: the ability to connect complex, descriptive text, such as “3 mm nodule in the left lower lobe”, to its precise anatomical location in three-dimensional space, a capability essential for grounded radiology report generation systems. The dataset comprises 3,142 non-contrast chest CT scans paired with standardized radiology reports from the CT-RATE dataset. Using a systematic three-stage pipeline, GPT-4 was used to extract positive lung and pleural findings, which were then manually segmented by expert annotators. A total of 8,028 findings across 16,301 entities were annotated, with quality control performed by board-certified radiologists. Approximately 79% of findings are focal abnormalities, while 21% are non-focal. The training set includes up to three representative segmentations per finding, while the validation and test sets contain exhaustive labels for each finding entity. ReXGroundingCT establishes a new benchmark for developing and evaluating sentence-level grounding and free-text medical segmentation models in chest CT. The dataset can be accessed at https://huggingface.co/datasets/rajpurkarlab/ReXGroundingCT.
@@ 1 Introduction
A central challenge for GRRG lies in bridging the gap between unstructured, free-text radiology reports and precise three-dimensional localization of findings within chest CT scans. Translating these narrative descriptions into spatially grounded labels remains largely unsolved. Existing chest CT datasets provide categorical labels or lesion masks but cannot represent the rich variability of free-text language [10, 20, 5].
We introduce ReXGroundingCT, the first publicly available manually annotated dataset connecting free-text radiology findings to precise 3D spatial localizations in chest CT imaging. Unlike previous data sets focused on predefined disease categories, ReXGroundingCT captures a diverse descriptive language and pairs it with detailed 3D segmentation masks. The dataset was developed via a multi-stage pipeline involving report standardization, finding extraction and categorization, and pixel-level annotation of CT scans by trained annotators and board-certified radiologists. An example of a data point from ReXGroundingCT is shown in Figure 1, A.
@@ 5 Discussion
Linking free-text radiology reports to precise 3D segmentations also has direct clinical relevance. Such tools could reduce the time and effort required for radiologists and clinicians to correlate narrative reports with imaging findings, facilitate clearer communication with referring physicians, and potentially reduce unnecessary consultations [14]. For patients, visual context directly linked to report language may improve comprehension and engagement with their health information [16]. Radiology trainees stand to benefit from systems that visually connect descriptive findings to anatomical regions, supporting the development of spatial reasoning skills essential for diagnostic accuracy [15].
@@ Appendix D Hierarchical Tree Finding Categories
[…]
.1 1) Typically non-focal lung/airway/pleural abnormalities. .2 1a) Bronchial wall thickening. .2 1b) Bronchiectasis. .2 1c) Emphysema (including Centrilobular, Paraseptal, Bullous). .2 1d) Septal thickening (including Interlobular, Reticulation). .2 1e) Micronodules (including Centrilobular, Tree-in-bud, Perilymphatic). .2 1f) Other. .1 2) Typically focal lung/airway/pleural opacities. .2 2a) Linear (including subsegmental atelectasis, scarring, fibrosis). .2 2b) Atelectasis, consolidation. .2 2c) Groundglass opacity. .2 2d) Pulmonary nodules/masses. .2 2e) Pleural effusion or thickening. .2 2f) Honeycombing. .2 2g) Pneumothorax. .2 2h) Other. .1 3) Non-pulmonary lesions. .2 3a) Lymphadenopathy lesions. .2 3b) Liver lesions. .2 3c) Gallbladder lesions. .2 3d) Renal/kidneys, collecting system, and ureters lesions. .2 3e) Spleen lesions. .2 3f) Adrenal lesions. .2 3g) Pancreas lesions. .2 3h) Thyroid lesions. .2 3i) Skin/subcutaneous lesions. .2 3j) Bone/Osseous structures lesions. .2 3k) Other lesions. .1 4) Bones (non-lesion). .2 4a) Fractures. .2 4b) Degenerative joint disease, degenerative disc disease, arthritis. .2 4c) Spinal curvature abnormalities: kyphosis, scoliosis. .2 4d) Other. .1 5) Stones/organ calcifications (non-lesion). .2 5a) Nephroliths, choleliths. .2 5b) Granulomas. .2 5c) Other. .1 6) Hollow viscera abnormalities. .2 6a) Hiatus hernia. .2 6b) Wall thickening. .2 6c) Dilated. .2 6d) Diverticulum (including diverticulosis). .2 6e) Other. .1 7) Skin/subcutaneous (non-lesion). .2 7a) Skin thickening. .2 7b) Stranding. .2 7c) Abdominal wall hernia: ventral, umbilical, inguinal. .2 7d) Gynecomastia. .2 7e) Other. .1 8) Cardiovascular. .2 8a) Atherosclerosis (including coronary, non-coronary). .2 8b) Vessel aneurysm, ectasia, enlargement. .2 8c) Vessel occlusion or stenosis. .2 8d) Cardiac chamber enlargement. .2 8e) Valvular calcification. .2 8f) Pericardial effusion. .2 8g) Other. .1 9) Body composition. .2 9a) Visceral fat. .2 9b) Superficial subcutaneous fat. .2 9c) Skeletal muscle. .2 9d) Osteoporosis. .2 9e) Hepatic steatosis. .2 9f) Other. .1 10) Diffuse/whole organ. .2 10a) Organomegaly (including splenomegaly, multinodular goiter, thyromegaly, lung hyperinflation). .2 10b) Atrophy. .2 10c) Other. .1 11) Device. .2 11a) Elongated (including catheter, pacemaker/defibrillator, spinal stimulator). .2 11b) Surgical clips. .2 11c) Other. .1 12) Other.

### v2

@@ Abstract
BACKGROUND Connecting free-text descriptions such as “3 mm nodule in the lower left lobe” to precise 3D segmentations remains an unsolved challenge in medical AI. Existing chest CT datasets rely on structured labels or predefined categories, limiting their ability to represent the richness of clinical language and support grounded radiology report generation systems. Bridging this gap requires datasets that capture the expressiveness of free-text radiology findings and link them to anatomically accurate spatial annotations in volumetric imaging.
METHODS We introduce ReXGroundingCT, the first publicly available dataset that links free-text findings to pixel-level 3D segmentations in chest CT scans. The dataset includes 3,142 non-contrast chest CT scans paired with standardized radiology reports from CT-RATE. Construction followed a structured three-stage pipeline. First, GPT-4 was used to extract and standardize findings, descriptors, and metadata from radiology reports originally written in Turkish and machine-translated into English. Second, GPT-4o-mini categorized each finding into a hierarchical ontology of lung and pleural abnormalities. Third, 3D annotations were produced on all CT volumes: the training set was annotated by quality assured by board-certified radiologists, while the validation and test sets were labeled completely by board-certified radiologists. Additionally, we constructed a complementary chain-of-thought dataset that provides step-by-step hierarchical anatomical reasoning for localizing findings within the CT volume, using GPT-4o and guided by localization coordinates derived from organ segmentation models.
RESULTS ReXGroundingCT provides 16,301 annotated entities across 8,028 text-to-3D-segmentation pairs spanning diverse radiological patterns from 3,142 non-contrast CT scans. Approximately 79% of findings are focal abnormalities, while 21% are non-focal. The dataset includes a public validation set of 50 cases and a private test set of 100 cases, both annotated by board-certified radiologists. Model performance on the private test set is hosted on a public leaderboard at https://rexrank.ai/ReXGroundingCT.
CONCLUSIONS ReXGroundingCT is the first manually curated dataset linking free-text chest CT findings to 3D segmentation masks, establishing a valuable benchmark for developing and evaluating free-text medical segmentation models. It sets the foundation for enabling free-text finding segmentation and grounded radiology report generation in CT imaging. The dataset can be accessed at https://huggingface.co/datasets/rajpurkarlab/ReXGroundingCT.
@@ 1 Introduction
A central challenge for GRRG lies in bridging the gap between unstructured, free‑text radiology reports and precise three‑dimensional localization of findings within chest CT scans. Translating these narrative descriptions into spatially grounded labels remains largely unsolved. Existing chest CT datasets provide categorical labels or lesion masks but cannot represent the rich variability of free‑text language [11, 22, 5].
We introduce ReXGroundingCT, the first publicly available manually annotated dataset connecting free‑text radiology findings to precise 3D spatial localizations in chest CT imaging. Unlike previous data sets focused on predefined disease categories, ReXGroundingCT captures a diverse descriptive language and pairs it with detailed 3D segmentation masks. The dataset was developed via a multi‑stage pipeline involving report standardization, finding extraction and categorization, and pixel‑level annotation of CT scans by trained annotators and board‑certified radiologists. An example of a data point from ReXGroundingCT is shown in Figure 1, A.
@@ 2.4 Annotation
2.5 Anatomical Chain-of-Thought
Radiologists naturally follow a systematic spatial reasoning process when interpreting chest CT scans, mentally navigating from broad anatomical regions to precise locations where findings manifest. To enable AI systems to replicate this clinical reasoning pattern, we constructed a chain-of-thought (CoT) dataset that captures the hierarchical anatomical thought process radiologists employ when localizing findings within 3D chest CT volumes.
Motivated by this observation, we built the CoT dataset to complement ReXGroundingCT by explicitly modeling how radiologists reason through anatomical hierarchy when identifying abnormalities. It provides step-by-step hierarchical reasoning sequences that describe how to localize each finding within a 3D chest CT volume. Each reasoning trace begins at the whole-CT level and proceeds through relevant anatomical structures (e.g., lung, lobe, segment) to the region containing the finding, integrating spatial cues such as adjacency to major vessels or pleural surfaces. GPT-4o was used to generate these CoT traces, guided by bounding box coordinates of organs and lobes derived from RadGenome-Chest CT [25]. These CoT annotations offer a structured view of spatial reasoning that can support multimodal alignment or serve as supervision for reasoning-augmented segmentation models. A representative CoT example from the dataset is shown in Figure 5.
To assess the quality of the CoT dataset, we performed human validation on a randomly selected subset of 150 findings. Each CoT was evaluated by a senior radiologist according to two criteria: (1) finding alignment, ensuring that no descriptive details were added or omitted relative to the original finding, and (2) accuracy of inferred adjacency, verifying that spatial inferences about neighboring anatomical structures (e.g., “adjacent to pleura,” “near fissure”) were anatomically correct. The validation showed a mean alignment accuracy of 0.92, indicating that the majority of generated reasoning traces were consistent with the original finding descriptions. For the adjacency inference, the accuracy of 0.75, indicating that most inferred spatial relationships were plausible but some CoTs incorrectly assumed pleural or chest wall adjacency when not stated. Only cases in which the model made an adjacency inference were evaluated for this criterion, which was only in around 20% of the total cases evaluated.
Figure 5: Example of an anatomical chain-of-thought (CoT) reasoning chain. The example illustrates a coarse-to-fine reasoning chain that progressively narrows localization from the whole CT volume to the relevant organ, lobe, and adjacent structures, producing a step-by-step description aligned with the finding text. The lung and lobe boundaries are generated automatically from organ segmentation models [25].
@@ 3 ReXGroundingCT
4 Results
We evaluated several state-of-the-art text-prompted 3D medical segmentation models, including BiomedParsev2 [27], SAT [29], and SegVol [12], on the ReXGroundingCT test set. As shown on the public leaderboard hosted at ReXRank [26], all existing models perform poorly in grounding free-text findings to their corresponding 3D regions, underscoring the difficulty of this task due to the absence of prior datasets that link narrative CT findings to 3D annotations. All off-the-shelf models struggled to localize free-text findings in 3D space, achieving near-zero performance across metrics.
We also show that fine-tuning on ReXGroundingCT substantially improved segmentation accuracy. The fine-tuned SAT model (SAT-FT) achieved a Global Dice of 0.209, Global HIT Rate of 0.473, representing a significant improvement over the baselines. This demonstrates that ReXGroundingCT provides the necessary supervision signal to adapt existing architectures to the challenging task of free-text 3D grounding. Yet, substantial progress still to be made in terms of model architecture, data scale, and representation learning to achieve clinically meaningful performance for this task.
To facilitate continued progress on this task, an active leaderboard is maintained through ReXRank, allowing the community to benchmark new models and track performance.
@@ 6 Discussion
Linking free-text radiology reports to precise 3D segmentations also has direct clinical relevance. Such tools could reduce the time and effort required for radiologists and clinicians to correlate narrative reports with imaging findings, facilitate clearer communication with referring physicians, and potentially reduce unnecessary consultations [16]. For patients, visual context directly linked to report language may improve comprehension and engagement with their health information [18]. Radiology trainees stand to benefit from systems that visually connect descriptive findings to anatomical regions, supporting the development of spatial reasoning skills essential for diagnostic accuracy [17]. These use cases and tasks are illustrated in Figure 6.
Figure 6: Potential downstream applications of ReXGroundingCT. Training on the ReXGroundingCT dataset can enable models for two key tasks: (1) report grounding, which localizes findings described in a new patient’s report, and (2) grounded report generation, which produces spatially grounded reports directly from CT scans. These capabilities support three key downstream use cases: (1) radiologist education, exemplified by RadGame [6], an AI-powered platform designed to improve localization and report-writing capabilities of future radiologists; (2) patient explanation, demonstrated by ReXplain [18], which visually communicates findings on patient scans to enhance understanding; and (3) report visualization, which augments radiology reports with spatially grounded imaging overlays to improve communication between radiologists and referring physicians.
@@ 8 Acknowledgement
This research was also supported by Harvard Medical School Dean’s Innovation Award for Accelerating Foundation Model Research
@@ Appendix D Hierarchical Tree Finding Categories
1) Typically non-focal lung/airway/pleural abnormalities
    1a) Bronchial wall thickening
    1b) Bronchiectasis
    1c) Emphysema (including Centrilobular, Paraseptal, Bullous)
    1d) Septal thickening (including Interlobular, Reticulation)
    1e) Micronodules (including Centrilobular, Tree-in-bud, Perilymphatic)
    1f) Other
    2) Typically focal lung/airway/pleural opacities
    2a) Linear (including subsegmental atelectasis, scarring, fibrosis)
    2b) Atelectasis, consolidation
    2c) Groundglass opacity
    2d) Pulmonary nodules/masses
    2e) Pleural effusion or thickening
    2f) Honeycombing
    2g) Pneumothorax
    2h) Other
    3) Non-pulmonary lesions
    3a) Lymphadenopathy lesions
    3b) Liver lesions
    3c) Gallbladder lesions
    3d) Renal/kidneys, collecting system, and ureters lesions
    3e) Spleen lesions
    3f) Adrenal lesions
    3g) Pancreas lesions
    3h) Thyroid lesions
    3i) Skin/subcutaneous lesions
    3j) Bone/Osseous structures lesions
    3k) Other lesions
    4) Bones (non-lesion)
    4a) Fractures
    4b) Degenerative joint disease, degenerative disc disease, arthritis
    4c) Spinal curvature abnormalities: kyphosis, scoliosis
    4d) Other
    5) Stones/organ calcifications (non-lesion)
    5a) Nephroliths, choleliths
    5b) Granulomas
    5c) Other
    6) Hollow viscera abnormalities
    6a) Hiatus hernia
    6b) Wall thickening
    6c) Dilated
    6d) Diverticulum (including diverticulosis)
    6e) Other
    7) Skin/subcutaneous (non-lesion)
    7a) Skin thickening
    7b) Stranding
    7c) Abdominal wall hernia: ventral, umbilical, inguinal
    7d) Gynecomastia
    7e) Other
    8) Cardiovascular
    8a) Atherosclerosis (including coronary, non-coronary)
    8b) Vessel aneurysm, ectasia, enlargement
    8c) Vessel occlusion or stenosis
    8d) Cardiac chamber enlargement
    8e) Valvular calcification
    8f) Pericardial effusion
    8g) Other
    9) Body composition
    9a) Visceral fat
    9b) Superficial subcutaneous fat
    9c) Skeletal muscle
    9d) Osteoporosis
    9e) Hepatic steatosis
    9f) Other
    10) Diffuse/whole organ
    10a) Organomegaly (including splenomegaly, multinodular goiter, thyromegaly, lung hyperinflation)
    10b) Atrophy
    10c) Other
    11) Device
    11a) Elongated (including catheter, pacemaker/defibrillator, spinal stimulator)
    11b) Surgical clips
    11c) Other
    12) Other